OpenStudy (anonymous):

Someone please: If ray PQ bisects angle SPF, and if the measurement of angle SPQ = 3x-2 and the measurement of angle MPF = 2x+9 then find the measurement of angle SPF

OpenStudy (turingtest):

do you mean angle QPF=2x+9? there is no MPF in the problem and do you want to know how to do it or just the answer?

OpenStudy (anonymous):

Yes, sorry, QPF.

OpenStudy (turingtest):

Answer or process?

OpenStudy (anonymous):

process

OpenStudy (turingtest):

ok let me draw a picture...

OpenStudy (anonymous):

Are you drawing?

OpenStudy (turingtest):

we are told PQ BISECTS angle SPF, which means it divides it into two congruent angles. so the measure of those angles are equal, i.e. SPQ=QPF 3x-2=2x+9 x=11 SPF=SPQ+QPF=3x-2+2x+9=5x+7 plug in x=11: SPF=5(11)+7=62

OpenStudy (anonymous):

Thank you, thank you.

OpenStudy (turingtest):

you're welcome